### v1

# AI sustains higher strategic tension than humans in chess

Abstract
Strategic decision-making involves managing the tension between immediate opportunities and long-term objectives. We study this trade-off in chess by characterizing and comparing dynamics between human vs. human and AI vs. AI games. We propose a network-based metric of piece-to-piece interaction to quantify the ongoing strategic tension on the board. Its evolution in games reveals that the most competitive AI players sustain higher levels of strategic tension for longer durations than elite human players. Cumulative tension varies with algorithmic complexity for AI and correspondingly in human-played games increases abruptly with expertise at about 1600 Elo and again at 2300 Elo. The profiles reveal different approaches. Highly competitive AI tolerates interconnected positions balanced between offensive and defensive tactics over long periods. Human play, in contrast, limits tension and game complexity, which may reflect cognitive limitations and adaptive strategies. The difference may have implications for AI usage in complex, strategic environments.
Chess — Artificial Intelligence — Complex Networks — Strategic Tension — Cognitive Science
preprint: APS/123-QED
The aphorism that one may have won the battle but lost the war is encapsulated in the notion of a “Pyrrhic victory.” Costly short-term wins must be balanced against the longer-term uncertainties, opportunities, or challenges that may emerge in competitive environments. This presents a fundamental challenge in strategic decision-making for both biological and artificial agents across diverse competitive and cooperative landscapes [42, 37, 3], of which war [43, 4] and conflicts [39, 11, 5] are examples. In the more general case, organisms and individuals learn or, over evolutionary time scales, adapt to changing ecological niches such as during development (e.g., neural pruning [44]), foraging (e.g., depending on climate [28]), and social change (e.g., reorganization of social hierarchy [21]). These can be characterized as strategic choices involving an implicit trade-off between overfitting the local environment and maintaining the flexibility to adjust to unforeseen environmental conditions [19]. A complete measurement of the strategic environment, costs, and even the outcome is not readily available in many cases. The game of chess, in contrast, presents a long-studied and bounded example of competitive strategic interaction, where the decisions made by the players and their consequences for the game outcome are completely measurable.
As a result, the game of chess has been the model system for investigating strategy. It has well-defined rules, perfect information, complexity, and rich history as a testbed for both human expertise and artificial intelligence (AI) [33, 38, 8, 29]. Chess is a competitive game between two white and black players who take turns making moves with a single piece on a board with 8x8 squares, as shown in Figure 1. Both players start with 16 pieces, and the game aims to capture the enemy’s king. A key feature that makes the game interesting is the combinatorial explosion in potential moves that prohibits exhaustive calculation, and human players must thus make heuristic decisions about the potential long-term value of moves [38]. The nature of the trade-off is shaped by cognitive limitations, psychological pressures, and time constraints, often leading to biased or suboptimal decisions [16, 17, 15]. In contrast, modern AI, particularly deep reinforcement learning systems like AlphaZero [35] and Leela Chess Zero [20], display superhuman performance by developing strategies through self-play. Their strategies differ significantly from human approaches [36, 34, 22]. Alongside these, widely used engines like Stockfish [40] rely primarily on brute-force search with handcrafted evaluation functions and deep search trees, incorporating machine learning only marginally. In principle, this suggests that the nature of the trade-offs that they need to make are distinct from human players, but there is little characterization of properties that distinguish their gameplay.
Here, we compare humans to AI play by analyzing the network structure of piece-to-piece interactions [12] over many chess games. We consider thousands of games across a broad range of scenarios, including games with players across all skill levels and from official AI tournaments between competitive chess engines such as Leela Chess Zero and Stockfish, held between 2019 and 2024. In these tournaments, the engines operate adaptively, allocating varying amounts of time to different moves. To these, we add games generated locally with limited computational power, in which Stockfish is run at fixed search depths and thus with constant computational effort per move.
We find that the complexity of the game, as measured by “strategic tension,” follows a regular curve, whose features correspond to the different stages of the game. The curve is characteristically different for human plays compared with AI plays. Our observations indicate that humans and AI behave differently in complex strategic environments and raise questions about the repercussions of AI managing competitive settings.
I Defining Tension in Chess
Given the state of a chessboard as in Figure 1a, we consider the set of threatening and defensive relationships at each ply (half-move). The threatening piece would be one able to capture an opposing piece in one move. The defending piece would threaten any piece that would capture the piece it is defending. This is the corresponding weighted network in Figure 1b, where a node is either a chess piece $p_{i}$ or a vacant square $s_{i}$, and the links can be of three types:
  • Attack link: We define an attack link from $p_{i}$ to $p_{j}$ if $p_{i}$, a piece from one side, can legally capture $p_{j}$, a piece from the opposing side. Attack links reflect direct threats, highlighting where immediate exchanges could occur.
  • Defense link: We define a defense link from $p_{i}$ to $p_{j}$ if there exists an opponent’s piece $p_{k}$ that can legally capture $p_{j}$ and after $p_{k}$’s move, $p_{i}$ would be able to legally capture $p_{k}$ in a single move. Kings can defend other pieces, but cannot be defended under this definition. Defense links reveal the layers of protection that influence whether an attack can be safely carried out or needs to be postponed.
  • Control link: We define a control link from $p_{i}$ to a vacant square $s_{j}$ if there exists an opponent’s piece $p_{k}$ that can legally move to $s_{j}$ and after $p_{k}$’s move, $p_{i}$ would be able to legally capture $p_{k}$ on $s_{j}$ in a single move. Control links, finally, represent subtler forms of influence, where empty squares become contested zones.
In constructing the network, we disregard which side is to move, focusing instead on the full set of interactions at each ply to capture the full state of the board. Attack and defense links can have a weight of 1 or 2, depending on whether the corresponding conditions are satisfied by one or both pieces involved. In contrast, control links always have a weight of 1. By construction, the network has at most 64 connected nodes, each corresponding to either a chess piece or a vacant square. For our work, we do not distinguish between link types  [^1] Although we also experimented with alternative networks that distinguish between link types and assign values according to the pieces involved, following commonly accepted conventions, these variations yielded qualitatively similar results in our studies. We therefore opted for the simpler representation to maintain interpretability and focus on the main structural features of the positions.. The resulting network captures the interplay among pieces and squares that creates and sustains tension on the chessboard.
Figure 1: One example of tension network. (a) Chess board at the ply 44 during a game between the two grandmasters Magnus Carlsen (white) and Alexey Sarana (black) in the Titled Tuesday tournament, Round 6, held on chess.com on November 26, 2024. (b) The network represents the interactions on the chessboard. The chess pieces are depicted as dark and light circles, depending on whether they are black or white. Attack links are orange, when originating from white pieces, and red, when originating from black. Defense links are turquoise when white is defending and blue otherwise. Control links are light blue (originating white) and light green (originating black). Vacant squares under control are shown as small brownish nodes. The chosen ply represents the highest tension reached in each game according to the maximum eigenvalue of the adjacency matrix.
Among the many ways of characterizing the interaction network, we are particularly interested in how the complexity of the game evolves, captured in features like the number of interacting pieces and loops  [^2] A loop consists of a closed sequence of interactions between pieces that form a cycle within the network (e.g., a white rook on b4 defends a white knight on e4; the knight, in turn, attacks a black bishop on d6, which is attacking the rook on b4). of interacting pieces, as described above. Importantly, we interpret this complexity not just in terms of structural richness, but also in terms of strategic fragility, the potential for cascading consequences following a suboptimal move. In dense, highly connected configurations, even a small error may propagate across the network, leading to substantial shifts in advantage. To capture this latent vulnerability, we use the maximum eigenvalue of the adjacency matrix, $\lambda_{1}$, for which we borrow an analogous, technical term from chess analysis, “strategic tension.” This quantity summarizes how tightly coupled the interaction network is at any given moment, and has been shown to reflect critical structural dynamics [27]. For simplicity, we consider the undirected tension network, which ensures a unique and real-valued largest eigenvalue.
The maximum eigenvalue reflects network functionality, including its dynamical properties and specific structural features [6]. This connection can be understood through the lens of topological entropy, a concept from dynamical systems that captures the complexity and unpredictability of paths in a network [1]. In this context, the spectral radius—the largest eigenvalue—reflects the network’s inherent “diversity” of possible trajectories. A higher value indicates greater dynamical complexity, mirroring the idea of tension as a measure of uncertainty or potential instability in the system. Importantly, the spectral radius is bounded between the average and the maximum node degree, i.e., $\langle s\rangle\leq\lambda_{1}\leq s_{\max}$, which emphasizes its role as a structural proxy capturing both global connectivity and local heterogeneity [41]. Thus, the spectral radius serves as a natural metric for tension, bridging structural properties (as ensured by the Perron–Frobenius theorem) and dynamical behavior (as reflected in entropy-like measures) as a measure of game complexity at the current state of the board (more details in the Supplementary Materials).
Given a network with $N=64$ nodes and an adjacency matrix whose entries are at most 2, the highest possible maximum eigenvalue is T=2(N1)=126T=2(N-1)=126, which occurs when the network is fully connected and each edge has weight 2. However, in the context of chess games, the adjacency matrices are quite sparse, and the resulting tension is significantly lower than this theoretical upper bound. Due to the constraints imposed by legal chess moves, such a fully connected weighted network is not achievable in practice. The example shown in Figure 1, which corresponds to the ply with the highest tension, has a tension value of approximately $T\simeq 6.65$. For reference, the maximum tension that can be reached by two players iteratively choosing legal moves that most increase tension is around $T\simeq 10.3$ at ply 52, with limited room for further increases (see Supplementary Materials).
Result 1: AIs tolerate more tension
We show the typical evolution of the strategic tension throughout many games in Figure 2. We can distinguish four main phases. The first phase, corresponding roughly to the opening stage up to around ply 30, features a buildup where pieces start approaching each other, and contested squares come under threat, leading to an abrupt increase in tension. The second phase, spanning approximately from ply 30 to ply 50, represents the peak of complexity, where the board remains highly interconnected and the tension stays elevated. The third exhibits a simplification of the tension network, as exchanges and reductions gradually lower the overall tension to about ply 120. The final and fourth is characterized by a low but relatively stable tension, as the remaining pieces maneuver without significantly altering the balance. The third and fourth phases are strongly influenced both by the overall game length and by the eventual outcome. Thus, the tension captures the overall contours of the game, as discussed in the theory of chess play [10], and indicates that the maximum eigenvalue is a meaningful proxy for game complexity and the potential for cascading effects. We do not report tension values beyond ply 150, as the available statistics become increasingly sparse—especially for games played between humans, which typically conclude earlier (Figure 2b).
We compare the tension profile with the Stockfish evaluation of the same games, which quantifies the game board’s state. We find that sharp changes in evaluation often occur at or shortly after peaks in tension [2], supporting the idea that these moments mark critical points where a single suboptimal move can trigger significant consequences (see Supplementary Materials).
Figure 2: Tension values during games for both humans (grandmasters) and AI (Stockfish and Leela Chess Zero). We consider 1,200 games for each, with human games obtained from https://www.pgnmentor.com and AI games from https://tcec-chess.com, which hosts official engine-versus-engine tournaments (a) Average tension is represented using cold and warm colors for humans and AI, respectively. For both groups, we distinguish between games that end in a draw and those that result in a win/loss, represented with light and dark shades, respectively. (b) For human games, the plots show the percentage of games—relative to the original total—that end in a draw or in a decisive result (win/loss) at each ply. Initially, the dataset is split between draws and decisive outcomes, and both percentages steadily decline as the number of plies increases, indicating how long games tend to last statistically. (c) The same analysis is performed for AI games, revealing a similar trend.
Following the peak, the rate at which the tension decreases depends on several factors. Among these, whether a game ends in a draw or not plays only a marginal role compared to the distinction between human-played and engine-played games. Generally, draws are significantly more common among chess engines, and tension lasts longer (Figure 2c). Moreover, the peak tension typically occurs at a later ply and tends to be lower on average compared to human games. Within human games, we observe that those ending in a draw tend to exhibit lower tension. This is likely due to many drawn games’ more cautious and symmetrical nature, where early simplifications and fewer sharp imbalances keep the tension low. In contrast, tension values are not as clearly differentiated between drawn and decisive outcomes in games played between AIs.
Result 2: Complementary Metrics Beyond Tension
Tension captures general and coarse features of the game. For example, AI maintains a higher number of pieces on average, reflecting less simplification early on. The interaction networks derived from engine games display more links and loops, indicating a richer and more resilient web of mutual threats, defenses, and controls. That the maximum average number of loops is, on average, higher in human games, as with the higher peak in tension, indicates similarity between the two metrics, which are discussed further in “ Supplementary Materials.”
Figure 3: Comparison of structural properties across human and AI games. We analyze the set of games shown in Figure 2. Two measures are considered: the average standard deviation of the degrees associated with chess pieces, normalized by the mean of the degrees, and the average tension per piece. Red is for AI and blue for human games.
As with tension, engine play in the endgame shows more variation in how pieces are connected to one another. We measure this with the noise coefficient of the degree distribution over chess pieces (standard deviation over the mean degree), which follows a similar trend in both human and AI games up until the final stages, where they start to diverge at approximately ply 80 (Figure 3a). There, AI exhibits greater variability in the degree distribution. This indicates that, on average, human players tend to distribute interactions more evenly across their pieces, while AI tends to stress interactions on specific pieces. This pattern aligns with the higher tension values in the endgame because a more heterogeneous distribution of degrees facilitates the emergence of highly connected hubs, increasing the adjacency matrix’s maximum eigenvalue.
This difference in tension between AI and human play, however, is inverted when normalized by the number of chess pieces remaining on the board, $\lambda_{1}/n_{t}$. As shown in Figure 3b, it remains consistently higher across all plies in human games. While human play shows an approximately linear increase in tension per piece throughout the game, AI games display a sublinear progression that appears to plateau around ply 80. This suggests that while AI maintains “tactical pressure” more persistently during late-game phases compared to human play, it does so with significantly more pieces remaining on the board. To quantify better this effect, we also investigated the correlation between tension values and the number of chess pieces present on the board (see Supplementary Materials).
While we have thus far treated edges in the graph equally, they are different. In particular, the surplus of attack links over defense links, or Attack-Defense Balance (Figure 4), highlights how the early phase of the game tends to be more defense-oriented. As the game progresses, however, attack links gradually outnumber defense links. This shift is noticeably more pronounced in human games compared to those played by AI.
Figure 4: Balance between attack and defense links across human and AI games. We analyze the set of games shown in Figure 2. For both human and AI games, we distinguish between games that end in a draw and those that result in a win/loss, represented with light and dark shades, respectively.
This behavior depends on different strategic styles, especially for drawn AI games, which a lower Attack-Defense Balance characterizes. Human players often seek direct and decisive operations, favoring sharper, more aggressive transitions when tactical chances arise. In contrast, AI maintains a more balanced configuration in the long run, resulting in wider piece coordination, latent threats, and structurally embedded defenses. In draws, AI games prefer solidity over risk, with extended defensive networks and fewer explicit attacks. This pattern also aligns with the tension trends shown in Figure 2, where tension remains high and decreases slowly in drawn AI games. Since defense links can only form in response to attacks, a high tension level in the later phases of the game likely reflects a dense but balanced interaction, where attacks tend to be answered by appropriate defenses. Extending the parallel with the tension profile, we observe that in this case, it is human games that reach a higher peak of tension, whereas in the case of the Attack-Defense Balance, AI games exhibit a more pronounced minimum. This comparison shows that tension aligns with a balance between attack and defense.
Overall, we find that strategic tension aligns with basic metrics about gameplay, such as the number of pieces and loops in the interaction graph. Still, it is likewise helpful to compare it with the balance across pieces on the board and how the pieces are arranged in formation. The picture is that AIs maintain more pieces, a more balanced network of attack and defense connections, and display more variability in the tension across the pieces they retain.
Result 3: Tension increases with skill
Figure 5: Tension loads for different Elo ratings and Stockfish depths. (a) We represent tension loads (sum of tension values up to ply 150) for human games (blue dots) and AI games between Stockfish and Leela from official TCEC tournaments (red dot), all associated with estimated Elo ratings. Each dot represents the average across 1,200 games. The blue dot associated with the highest Elo rating and the red dot were derived from the data analyzed in Figure 2. (b) We represent tension loads from games between versions of Stockfish at varying depth levels (orange dots), computed from a smaller sample of 120 games per dot.
While engines can maintain higher tension for a longer duration compared to humans, how does this difference depend on skill level? We consider the cumulative tension up to ply 150 across different Elo ratings for human players in Figure 5a  [^3] In principle, Stockfish has a search depth parameter that serves as a proxy for skill level. Leela has a time limit parameter, during which it is permitted to assess the state of the board and make a decision. At a few dozen search depths, the tension load is limited to about 450, similarly for Leela at a few seconds. Since the games that are shown go far beyond these limited parameter values with supercomputers, this suggests that algorithmic complexity increases the ability to handle tension.. Tension changes substantially with Elo, revealing discrete jumps that separate human players into roughly three, significantly separate groups. A plateau in tension load from Elo of 1000 ends between 1500 and 1700. This transition typically corresponds to the entry point into competitive chess, where players begin to study openings, apply strategic principles more consistently, and reduce the frequency of major tactical blunders. After a second plateau, a jump in tension occurs at 2100 Elo. This point marks the entry into the realm of stronger players, such as experts and grandmasters, whose games exhibit more nuanced, complex positions with deeper strategic and tactical layers.
Figure 5b shows the tension load values for various fixed search depths of Stockfish. These results indicate an approximately constant tension load across different depths, suggesting a qualitatively different scenario compared to the Stockfish vs. Leela games from the TCEC tournaments (highlighted as the red point in Figure 5a), where the engines are free to adjust their search depth and, more generally, the time allocated per move  [^4] Representing the dependence of tension on the time limit per move in the case of Leela is challenging because it strongly depends on the technical specifications of the machine used, particularly the presence of a suitable dedicated GPU and sufficient RAM. In particular, achieving Elo ratings significantly higher than the best human players requires substantial hardware resources.. The correlation between Stockfish search depth and Elo rating is not well established. Previous experiments on a different engine (Houdini) reported a roughly linear relationship up to a search depth of 20 plies [13].
II Discussion
We investigate the dynamics of strategic tension, a network-based definition of piece interactions, in chess games played by humans and AI. Our central hypothesis is that distinct tension profiles would emerge, reflecting differing strategic approaches. The results confirm this hypothesis and reveal several key insights. First, we observed that state-of-the-art chess engines (Stockfish at high search depths and Leela Chess Zero at advanced neural network settings) consistently maintain significantly higher levels of strategic tension for longer durations than even the best grandmasters. This suggests an ability to navigate and sustain complex, balanced positions where multiple pieces exert pressure and threats remain unresolved. Second, human players exhibit a notable relationship between skill level (Elo rating) and tension management. Significant differentiation in accumulated tension emerges primarily above an Elo rating of approximately 1600, coinciding with the stage where players typically develop deeper strategic understanding, theoretical knowledge, and enhanced calculation abilities [9, 7]. This implies that managing higher levels of strategic tension is characteristic of advancing chess expertise in humans. Likewise, one might expect greater computational complexity to allow an engine to find and sustain even more complex positions, leading to a monotonic increase in tension load with search depth, similar to the trend observed with human expertise. Preliminary calculations indicate that the computational complexity of Stockfish, as reflected in its search depth, does influence tension load, but in a nonlinear manner: tension load exhibits flat plateaus across a wide range of fixed search depths (approximately from 10 to 30). This raises the question of how and where new play styles emerge as the complexity accessible to the algorithm increases.
The observed differences in tension dynamics distinguish how biological and artificial systems approach problem-solving under strategic constraints. Human cognition appears adapted to manage, and perhaps even actively reduce, high levels of sustained complexity, possibly due to inherent limitations in working memory, attention, or susceptibility to psychological pressure [15, 18]. State-of-the-art AI, optimized for performance within defined rule sets, demonstrates a capacity for tolerating and exploiting complexity [30]. The superior ability of high-level engines to sustain tension elucidates their ability to handle immense combinatorial complexity [34, 22]: engines do not simply calculate better but actively maintain more complex strategic states. This aligns with the idea that optimal strategy might be contingent on the agent’s cognitive or computational architecture. More broadly, such differences may be essential for AI systems designed for human interaction. While direct extrapolation requires caution, we might think of how to examine tension in other complex adaptive systems, such as economic markets or geopolitical negotiations [32]. The finding that increased computational complexity allows AI to sustain higher tension may matter for stability in international relations, potentially enabling longer standoffs but also increasing the stakes should tension break [14, 31] in an AI-enabled analogy to Dr. Strangelove.
Several avenues for future research emerge from this work. The network-based tension metric represents one specific operationalization of strategic complexity. It primarily captures piece interactions and may not fully encompass other crucial factors like pawn structure, king safety, or long-term positional plans. While it changes predictably with several measures of board complexity (Result 2), parallel measures may help inform pedagogical approaches in chess. Further investigations could also focus on how different engines manage tension, examining variations across engine architectures, versions, and underlying algorithms. Extending this methodology to other complex strategic games, such as Go [36], Shogi, or even real-time strategy e-sports, could test the universality of the observed dynamics across different game structures (e.g., imperfect information, different spatial dynamics). Directly investigating the cognitive and physiological correlates of our tension metric in human players—for example, by combining game analysis with eye-tracking, think-aloud protocols, or stress indicators (like galvanic skin response)—could help elucidate the underlying mechanisms driving human tension. Furthermore, exploring whether AI can be explicitly trained to modulate a tension profile, perhaps to mimic human playstyles or achieve specific pedagogical goals, represents an interesting direction for human-AI interaction research. Investigating the relationship between our tension metric and game outcomes (win/loss/draw) beyond simple averages could also reveal critical thresholds or patterns predictive of game resolution. Finally, our analysis relies on existing game databases. While large, these datasets may have inherent biases, and game conditions (e.g., time controls, stakes) might differ from controlled experimental settings. The specific AI engine used (Stockfish) represents one architecture; other engines with different algorithms (e.g., neural network-based like AlphaZero beyond its initial papers) might exhibit different tension dynamics. While this study focused on Stockfish and Leela due to their prominence and availability, the analytical framework and open-source code we provide can be readily applied to other UCI-compatible engines, paving the way for broader comparative studies.
We find evidence for distinct approaches to managing strategic tension between human players and AI chess engines. By proposing a network-based metric, we characterize how tension typically accumulates and dissipates throughout games, revealing that high-performance AI sustains significantly greater complexity for longer periods than even elite human players. Human tension management appears correlated with expertise but potentially constrained by cognitive factors, while AI’s capacity in tension management is primarily a function of computational resources, which allows it to consider high-dimensional configuration of balanced interactions. While human players would find such algorithmic complexity inaccessible, the brain is the outcome of a trade-off in terms of longer-term evolutionary games. The short-term cost of losing a chess game might be worth the long-term benefits of metabolic cost and cognitive plasticity in the game of life.
Code availability
Python code used to compute tension and generate the Stockfish self-play games is available on GitHub at https://github.com/Draghellon/Tension_Chess.
Authors’ contributions
AC developed the code, performed the numerical analyses, and generated the figures. VDPS provided the original inspiration for the project and supervised its development at all stages. EDL participated in discussion. All authors contributed to writing the manuscript.
Acknowledgements
We thank V. Latora for valuable discussions, particularly regarding the inclusion of Fig. 4, and M. Barthelemy for useful discussions and for sharing his code, which we used as a baseline for our analysis. This work was partly funded by the City of Vienna, Municipal Department 7, and the Federal Ministry of the Republic of Austria for Climate Action, Environment, Energy, Mobility, Innovation, and Technology as part of the project GZ2021-0.664.668. It was also partially funded by the Austrian Science Fund (FWF) 10.55776/ESP127.
III Methods
III.1 Data
Chess games are commonly stored in the PGN (Portable Game Notation) format. This standard plain-text representation records the sequence of moves in a game, along with optional metadata such as player names, event, result, and time controls. This format is human-readable and machine-parsable, making it ideal for large-scale game analysis.
The games between top-level human players were obtained from https://www.pgnmentor.com, which provides thousands of PGN-formatted games organized by player or sorted by opening. This source includes a comprehensive archive of historical and modern games, making it well-suited for comparative analysis.
Competitive matches between Stockfish and Leela Chess Zero were collected from the https://tcec-chess.com/ website, which regularly publishes high-level engine-versus-engine games in PGN format. These games represent the cutting edge of chess engine performance under standardized tournament conditions. The typical time control is 120 minutes per player, with a 12-second increment added after each move.
These datasets served as the basis for the analyses presented in Figure 2.
In addition, a large set of games played by human players with Elo ratings ranging from 1100 to 2300 was downloaded from https://www.ficsgames.org/, an online archive that collects games played on the Free Internet Chess Server (FICS). This dataset broadly represents amateur to intermediate-level players, allowing for analysis across a wide skill spectrum.
Together with the previously described datasets, these data formed the basis for the analyses presented in Figure 5.
III.2 Software
All operations related to chess move generation, rule validation, and position analysis were carried out using the python-chess library (v1.9.4). This comprehensive library offers tools for representing chess boards, parsing PGN files, and computing legal moves according to FIDE rules. Furthermore, it provides native support for communicating with UCI-compatible engines such as Stockfish, which allowed us to automate the evaluation of positions and simulate engine-vs-engine games with varying search depths.
III.3 Generating Stockfish games at varying depths
To generate the games between Stockfish and itself at various search depths, we followed a procedure designed to balance realism and controlled variability. While there is no universally accepted mapping between Stockfish’s search depth and Elo rating, approximate correspondences have been estimated in prior works [13]. Based on these estimates, we selected a range of depths roughly aligned with the playing strength of top human players and beyond.
To introduce diversity in the opening phase and avoid deterministic repetition, each game was initialized using the first 12 plies (typically covering the opening) from randomly selected games played by top-level human players. After this opening phase, the rest of the game was played entirely by two instances of Stockfish, one playing each move at either depth $D$ or $D+1$, chosen randomly, and the other at either depth $D$ or D1D-1, also chosen randomly. This asymmetry introduced slight variability in evaluation and decision-making while preserving a consistent overall depth level $D$.
IV Supplementary materials
IV.1 Why the spectral radius as a measure of tension
Topological entropy, $h_{\mathrm{top}}$, is a concept from dynamical systems theory that quantifies the exponential growth rate of the number of distinguishable orbits or, in the context of graphs, the diversity of paths. For unweighted networks (where the adjacency matrix has entries in $\{0,1\}$), the spectral radius $\lambda_{1}$ of the adjacency matrix is directly related to topological entropy. In particular, it has been shown that $\lambda_{1}$ satisfies a variational principle equivalent to the dynamical entropy of the network (i.e., the entropy of a stochastic process on the graph), leading to the definition of topological entropy as
$h_{\mathrm{top}}=\log(\lambda_{1}).$
Topological entropy was first introduced in 1965 by Adler, Konheim, and McAndrew [1] as a measure of the complexity of a dynamical system. In the context of networks, this relationship is particularly meaningful: the quantity $\lambda_{1}$ does not merely reflect the number of walks in the graph, but its logarithm measures the average Shannon information per step gained when observing a random walk unfold. A larger $\lambda_{1}$ indicates a greater number of possible distinct paths at each step, implying higher unpredictability, richer structural diversity, and overall complexity.
This connection positions $\lambda_{1}$ as a fundamental quantity bridging structural graph theory, information theory, and dynamics. It can be interpreted as a proxy for the “intrinsic computational capacity” or “information-processing potential” of the network.
In our study, this interpretation is particularly useful: we use the spectral radius $\lambda_{1}$ as a global measure of the tension and complexity present on the chessboard. A higher $\lambda_{1}$ suggests a denser and more interconnected pattern of potential interactions between pieces (e.g., threats and defenses), and therefore a richer position in tactical possibilities and strategic depth. This justifies using $\lambda_{1}$ as a meaningful descriptor of the game’s complexity at a given state.
IV.2 Maximized tension game
If we consider the unrealistic and absurd game in which, at each ply, the players randomly choose one of the moves that maximizes tension, by ply 52 the tension reaches a value of $T\simeq 10.3$ (Figure 6).
Figure 6: Tension network obtained by maximizing tension move by move. (a) Chessboard at ply 44, obtained by choosing the legal move that increases the tension the most at each ply. (b) The corresponding network represents the interactions on the chessboard, using the same color and edge scheme as in Figure 1.
The first 60 moves of the game are listed below: 1. d3 h6 2. Bg5 e5 3. Be7 Nc6 4. Qd2 d5 5. Qxh6 Qd6 6. Qe6 Nd8 7. c4 Rh6 8. cxd5 e4 9. Nc3 g5 10. Na4 b5 11. Rc1 Rb8 12. Rxc7 Rb7 13. Nc5 Qxd5 14. dxe4 Rb6 15. e3 Ra6 16. Rxa7 b4 17. Bd3 Bg7 18. Nf3 Bf8 19. Nxg5 Bd7 20. f4 Bc8 21. Kf2 b3 22. Rc1 bxa2 23. Rc3 a1=Q 24. Ra3 Qd1 25. Ra4 Qg4 26. Be2 Qgf5 27. Bd3 Qfxe4 28. Ra8 Nxe7 29. Kg3 Rh4 30. Kf2 Rh6.
After ply 52, the game maintains a relatively constant level of tension indefinitely, indicating that the game ceases to progress meaningfully.
IV.3 Stockfish evaluation for humans and engines
To better analyze the progression of games, we also examined the Stockfish evaluation, again distinguishing between games played by humans or AIs, and whether they ended in a draw or not (Figure 7). What emerges is that games between AIs show a rapid increase in the absolute value of the Stockfish evaluation from the very first plies. As the game progresses, the average evaluation values begin to diverge depending on the eventual outcome: games that end in a win or loss exhibit a steady increase in evaluation with each ply, whereas those that result in a draw maintain relatively low evaluation values throughout.
In contrast, in human games, the evaluation remains near zero during the opening phase. It is worth noting that, in fact, only starting from 2023 did official rules begin allowing Leela Chess Zero to select its openings directly from a database. However, as the game progresses, a similar divergence in evaluation can be observed. Notably, this shift typically occurs around ply 40, where tension generally peaks, marking a transition where evaluations for decisive games begin to rise consistently. In contrast, those for drawn games remain comparatively stable. Barthelemy [2] has already observed this result, using a similar framework for capturing the notion of fragility within chess games.
In the case of human games, the average values of tension are characterized by greater fluctuations, likely due to the lower predictability of human play and, more specifically, the greater variability between individual games, especially when different players are involved. Moreover, when focusing on tension values associated with drawn games, a notable difference emerges: in the case of AI matches, tension decreases towards zero much more gradually. This is mainly because engine games are longer on average than human games.
Figure 7: Stockfish evaluation during games for humans and AI. We consider 120 games for each (a subset of those analyzed in Figure 2), with evaluations computed at a fixed depth of 20. The same color scheme described in Figure 2 is used: cold and warm colors for humans and AI, respectively, and light/dark shades to distinguish draws from decisive games. We also show the overall average evaluations, along with vertical dashed lines indicating the ply corresponding to the maximum tension for both groups, humans and AI, shown in blue and red, respectively.
To better understand the type of information conveyed by the Stockfish evaluation, we analyzed the 2,400 games to compute the probability that a player goes on to win the game when holding an evaluation advantage of at least 100 centipawns at ply 40. In the case of human games, this probability is approximately 71 %, whereas for AI games, it is around 59 %.
IV.4 Additional properties of the chessboard and network
In this section, we present further comparative metrics between human and AI chess games, including the number of pieces, the sum of the weighted material in the chess board - where pieces are valued at 1 (pawns), 3 (knights and bishops), 5 (rooks), 9 (queens), and 0 (kings) - along with the number of links and the number of loops in the tension network. These measures provide complementary insights into positional advantages and dynamic board conflicts across different types of games (Figure 8).
IV.5 Correlation between tension and the number of chess pieces
We investigated the correlation between the measured tension values and the number of chess pieces present on the board. Naturally, this correlation is highly dependent on the stage of the game being considered. For this reason, we analyzed three distinct moments in the game—specifically at ply 30, 60, and 90—to show how, on average, tension varies with the number of pieces on the board, both for the AI and for human players (Figure 9).
We observe that, on average, tension increases with the number of chess pieces across all three stages of the game. However, this increase diminishes rapidly as the number of pieces grows, eventually reaching an approximately constant behavior—or even a slight decrease. This may be because when the number of pieces is significantly higher than what would typically be expected at a given ply, it might indicate that no major confrontation has yet occurred. Such confrontations often lead to a rise in tension, as more pieces are brought into key positions on the board. Moreover, an excessive number of pieces on the board can severely limit their mobility, reducing their ability to control important areas or pose threats to the opponent. Therefore, it is reasonable to expect that, to maximize tension, there exists an ideal range for the number of pieces — neither too high nor too low.
Figure 8: Comparison of additional structural properties across human and AI games. We analyze the set of games shown in Figure 2. Four measures are considered: the number of pieces, the weighted material, the number of links, and the number of loops in the interaction network. The color scheme is red and blue for AI and human games, respectively.
Figure 9: Analysis of the dependence of tension on the number of chess pieces at three game moments. We analyze the set of games shown in Figure 2. Average tension values are shown for each number of chess pieces, distinguishing between three game moments: ply 30, 60, and 90. The color scheme is red and blue for AI and human games, respectively. The error bars correspond to the standard deviation calculated for each represented value.

### v2

@@
 Abstract
-Strategic decision-making involves managing the tension between immediate opportunities and long-term objectives.
-We study this trade-off in chess by characterizing and comparing dynamics between human vs.
-human and AI vs.
-AI games.
-We propose a network-based metric of piece-to-piece interaction to quantify the ongoing strategic tension on the board.
-Its evolution in games reveals that the most competitive AI players sustain higher levels of strategic tension for longer durations than elite human players.
-Cumulative tension varies with algorithmic complexity for AI and correspondingly in human-played games increases abruptly with expertise at about 1600 Elo and again at 2300 Elo.
-The profiles reveal different approaches.
-Highly competitive AI tolerates interconnected positions balanced between offensive and defensive tactics over long periods.
-Human play, in contrast, limits tension and game complexity, which may reflect cognitive limitations and adaptive strategies.
-The difference may have implications for AI usage in complex, strategic environments.
-Chess — Artificial Intelligence — Complex Networks — Strategic Tension — Cognitive Science
+Strategic decision-making requires balancing immediate opportunities against long-term objectives: a tension fundamental to competitive environments.
+We investigate this trade-off in chess by analyzing the dynamics of human and AI gameplay through a network-based metric that quantifies piece-to-piece interactions.
+Our analysis reveals that elite AI players sustain substantially higher levels of strategic tension for longer durations than top human grandmasters.
+We find that cumulative tension scales with algorithmic complexity in AI systems and increases linearly with skill level (Elo rating) in human play.
+Longer time controls are associated with higher tension in human games, reflecting the additional strategic complexity players can manage with more thinking time.
+The temporal profiles reveal contrasting approaches: highly competitive AI systems tolerate densely interconnected positions that balance offensive and defensive tactics over extended periods, while human players systematically limit tension and game complexity.
+These differences have broader implications for understanding how artificial and biological systems navigate complex strategic environments and for the deployment of AI in high-stakes competitive scenarios.
+Chess — Artificial Intelligence — Complex Networks — Strategic Tension
 preprint: APS/123-QED
-The aphorism that one may have won the battle but lost the war is encapsulated in the notion of a “Pyrrhic victory.” Costly short-term wins must be balanced against the longer-term uncertainties, opportunities, or challenges that may emerge in competitive environments.
-This presents a fundamental challenge in strategic decision-making for both biological and artificial agents across diverse competitive and cooperative landscapes [42, 37, 3], of which war [43, 4] and conflicts [39, 11, 5] are examples.
-In the more general case, organisms and individuals learn or, over evolutionary time scales, adapt to changing ecological niches such as during development (e.g., neural pruning [44]), foraging (e.g., depending on climate [28]), and social change (e.g., reorganization of social hierarchy [21]).
-These can be characterized as strategic choices involving an implicit trade-off between overfitting the local environment and maintaining the flexibility to adjust to unforeseen environmental conditions [19].
-A complete measurement of the strategic environment, costs, and even the outcome is not readily available in many cases.
-The game of chess, in contrast, presents a long-studied and bounded example of competitive strategic interaction, where the decisions made by the players and their consequences for the game outcome are completely measurable.
-As a result, the game of chess has been the model system for investigating strategy.
+The aphorism that one may have won the battle but lost the war is encapsulated in the notion of a “Pyrrhic victory.” Costly short-term gains must be balanced against the longer-term uncertainties, opportunities, or challenges that may emerge in competitive environments.
+This presents a fundamental challenge in strategic decision-making for both biological and artificial agents across diverse competitive and cooperative landscapes [42, 40, 3], of which war [weinsteinRebellionPolitics2007, bohorquezCommonEcology2009] and conflicts [smithLogicAnimal1973, dedeoInductiveGame2010a, brushConflictsInterest2018] are poignant examples.
+In the more general case, organisms and individuals learn such as during development (e.g., neural pruning [woudeNoGains2019]), foraging (e.g., depending on climate [piteraDailyForaging2018]), and social change (e.g., reorganization of social hierarchy [28]) or, over evolutionary time scales, “adapt” to changing ecological niches [24].
+In the context of an individual’s lifetime, these can be characterized as strategic choices that involve an implicit trade-off between overfitting the local environment and maintaining the flexibility to adapt to unforeseen environmental conditions [27], in an echo of the speed-accuracy trade-off [33].
+In many real-world settings, a complete measurement of the strategic environment, costs, and outcomes is not available.
+Chess, by contrast, is a game of perfect information, providing a well-studied and bounded setting.
+As a result, chess has served as a model system for investigating strategy.
 It has well-defined rules, perfect information, complexity, and rich history as a testbed for both human expertise and artificial intelligence (AI) [36, 39, 11, 32].
-Chess is a competitive game between two white and black players who take turns making moves with a single piece on a board with 8x8 squares, as shown in Figure 1.
-Both players start with 16 pieces, and the game aims to capture the enemy’s king.
-A key feature that makes the game interesting is the combinatorial explosion in potential moves that prohibits exhaustive calculation, and human players must thus make heuristic decisions about the potential long-term value of moves [38].
+Chess is a competitive game between two players, white and black, who take turns making moves with a single piece on a board with 8x8 squares as shown in Fig.
+1.
+Both players start with 16 pieces, and the objective of the game is to capture the enemy’s king.
+A key feature that makes the game interesting is the combinatorial explosion in potential moves, which prohibits exhaustive calculation.
+As a result, human players must make heuristic decisions about the possible long-term value of moves [39].
 The nature of the trade-off is shaped by cognitive limitations, psychological pressures, and time constraints, often leading to biased or suboptimal decisions [21, 25, 20].
-In contrast, modern AI, particularly deep reinforcement learning systems like AlphaZero [35] and Leela Chess Zero [20], display superhuman performance by developing strategies through self-play.
-Their strategies differ significantly from human approaches [36, 34, 22].
-Alongside these, widely used engines like Stockfish [40] rely primarily on brute-force search with handcrafted evaluation functions and deep search trees, incorporating machine learning only marginally.
-In principle, this suggests that the nature of the trade-offs that they need to make are distinct from human players, but there is little characterization of properties that distinguish their gameplay.
+In contrast, modern AI exhibits superhuman performance through two main strategies.
+One class involves deep reinforcement learning systems, such as AlphaZero and Leela Chess Zero, which develop strategies through self-play using deep convolutional neural networks.
+Alongside these, engines like Stockfish combine massive brute-force search trees with NNUE (Efficiently Updatable Neural Networks).
+While Stockfish originally relied on handcrafted evaluation functions, modern versions utilize machine learning to train their evaluation layers, blending classical search efficiency with neural network-based intuition.
+These differences suggest that the nature of the trade-offs that AI engines make is distinct from those made by human players.
+Still, there is little characterization of the properties that distinguish gameplay.
 Here, we compare humans to AI play by analyzing the network structure of piece-to-piece interactions [16] over many chess games.
-We consider thousands of games across a broad range of scenarios, including games with players across all skill levels and from official AI tournaments between competitive chess engines such as Leela Chess Zero and Stockfish, held between 2019 and 2024.
+We consider thousands of games across a broad range of scenarios, including games with players of all skill levels and official AI tournaments between competitive chess engines, such as Leela Chess Zero and Stockfish, held between 2019 and 2024.
 In these tournaments, the engines operate adaptively, allocating varying amounts of time to different moves.
-To these, we add games generated locally with limited computational power, in which Stockfish is run at fixed search depths and thus with constant computational effort per move.
-We find that the complexity of the game, as measured by “strategic tension,” follows a regular curve, whose features correspond to the different stages of the game.
-The curve is characteristically different for human plays compared with AI plays.
-Our observations indicate that humans and AI behave differently in complex strategic environments and raise questions about the repercussions of AI managing competitive settings.
+In addition to these tournament games, we include simulations in which Stockfish is run at fixed search depths.
+This serves as a controlled baseline, allowing us to evaluate piece-piece interaction networks under uniform computational constraints.
+We find that the complexity of the game, as measured by “strategic tension,” follows, on average, a regular curve, whose features correspond to the different stages of the game.
+The curve is characteristically different for human players compared with AI players.
+Our observations suggest that humans and AI exhibit distinct behaviors in complex strategic environments, raising questions about the consequences of delegating competitive real-life scenarios to AI.
 I Defining Tension in Chess
-Given the state of a chessboard as in Figure 1a, we consider the set of threatening and defensive relationships at each ply (half-move).
-The threatening piece would be one able to capture an opposing piece in one move.
+In chess, tension refers to a state of unresolved tactical interaction where multiple pieces from both sides occupy squares from which they can capture or be captured by one another.
+Unlike a direct exchange of pieces, which resolves the interaction and simplifies the board state, tension represents a “latent” complexity.
+It persists as long as players maintain these mutual threats without acting upon them, forcing both competitors to constantly evaluate the shifting consequences of a potential explosion of captures.
+By formalizing these threats into a network of offensive and defensive relationships, we can quantify this tension as the density and structure of piece-to-piece interactions.
+Given the state of a chessboard as in Fig.
+1a, we consider the set of offensive and defensive relationships at each ply (half-move).
+An attacking piece can capture an opposing piece in a single move.
 The defending piece would threaten any piece that would capture the piece it is defending.
-This is the corresponding weighted network in Figure 1b, where a node is either a chess piece $p_{i}$ or a vacant square $s_{i}$, and the links can be of three types:
+This is the corresponding weighted network in Fig.
+1b, where a node is either a chess piece $p_{i}$ or a vacant square $s_{i}$, and the links can be of three types:
   • Attack link: We define an attack link from $p_{i}$ to $p_{j}$ if $p_{i}$, a piece from one side, can legally capture $p_{j}$, a piece from the opposing side.
@@
 Kings can defend other pieces, but cannot be defended under this definition.
+This distinction arises because the King is never captured in chess; hence, the legal mechanism of a reciprocal capture (or “re-capture”) cannot apply to it.
 Defense links reveal the layers of protection that influence whether an attack can be safely carried out or needs to be postponed.
@@
 In constructing the network, we disregard which side is to move, focusing instead on the full set of interactions at each ply to capture the full state of the board.
-Attack and defense links can have a weight of 1 or 2, depending on whether the corresponding conditions are satisfied by one or both pieces involved.
-In contrast, control links always have a weight of 1.
+Attack and defense links can have weight 1 or 2, depending on whether the corresponding conditions are satisfied by one or both of the pieces involved.
+In contrast, control links always have weight 1.
 By construction, the network has at most 64 connected nodes, each corresponding to either a chess piece or a vacant square.
-For our work, we do not distinguish between link types  [^1] Although we also experimented with alternative networks that distinguish between link types and assign values according to the pieces involved, following commonly accepted conventions, these variations yielded qualitatively similar results in our studies.
+For our work, we do not distinguish between link types  [^1] Although we also experimented with alternative networks that distinguish between link types and assign values according to the pieces involved, following commonly accepted conventions, these variations yielded qualitatively similar results in our case studies.
 We therefore opted for the simpler representation to maintain interpretability and focus on the main structural features of the positions..
@@
 Figure 1: One example of tension network.
-(a) Chess board at the ply 44 during a game between the two grandmasters Magnus Carlsen (white) and Alexey Sarana (black) in the Titled Tuesday tournament, Round 6, held on chess.com on November 26, 2024.
+(a) Chessboard at the ply 44 during a game between the two grandmasters, Magnus Carlsen (white) and Alexey Sarana (black), in the Titled Tuesday tournament, Round 6, held on chess.com on November 26, 2024.
 (b) The network represents the interactions on the chessboard.
 The chess pieces are depicted as dark and light circles, depending on whether they are black or white.
-Attack links are orange, when originating from white pieces, and red, when originating from black.
+Attack links are displayed in orange when originating from white pieces and in red when originating from black.
 Defense links are turquoise when white is defending and blue otherwise.
 Control links are light blue (originating white) and light green (originating black).
-Vacant squares under control are shown as small brownish nodes.
-The chosen ply represents the highest tension reached in each game according to the maximum eigenvalue of the adjacency matrix.
-Among the many ways of characterizing the interaction network, we are particularly interested in how the complexity of the game evolves, captured in features like the number of interacting pieces and loops  [^2] A loop consists of a closed sequence of interactions between pieces that form a cycle within the network (e.g., a white rook on b4 defends a white knight on e4; the knight, in turn, attacks a black bishop on d6, which is attacking the rook on b4).
-of interacting pieces, as described above.
-Importantly, we interpret this complexity not just in terms of structural richness, but also in terms of strategic fragility, the potential for cascading consequences following a suboptimal move.
-In dense, highly connected configurations, even a small error may propagate across the network, leading to substantial shifts in advantage.
-To capture this latent vulnerability, we use the maximum eigenvalue of the adjacency matrix, $\lambda_{1}$, for which we borrow an analogous, technical term from chess analysis, “strategic tension.” This quantity summarizes how tightly coupled the interaction network is at any given moment, and has been shown to reflect critical structural dynamics [27].
-For simplicity, we consider the undirected tension network, which ensures a unique and real-valued largest eigenvalue.
-The maximum eigenvalue reflects network functionality, including its dynamical properties and specific structural features [6].
-This connection can be understood through the lens of topological entropy, a concept from dynamical systems that captures the complexity and unpredictability of paths in a network [1].
-In this context, the spectral radius—the largest eigenvalue—reflects the network’s inherent “diversity” of possible trajectories.
-A higher value indicates greater dynamical complexity, mirroring the idea of tension as a measure of uncertainty or potential instability in the system.
-Importantly, the spectral radius is bounded between the average and the maximum node degree, i.e., $\langle s\rangle\leq\lambda_{1}\leq s_{\max}$, which emphasizes its role as a structural proxy capturing both global connectivity and local heterogeneity [41].
-Thus, the spectral radius serves as a natural metric for tension, bridging structural properties (as ensured by the Perron–Frobenius theorem) and dynamical behavior (as reflected in entropy-like measures) as a measure of game complexity at the current state of the board (more details in the Supplementary Materials).
-Given a network with $N=64$ nodes and an adjacency matrix whose entries are at most 2, the highest possible maximum eigenvalue is T=2(N1)=126T=2(N-1)=126, which occurs when the network is fully connected and each edge has weight 2.
-However, in the context of chess games, the adjacency matrices are quite sparse, and the resulting tension is significantly lower than this theoretical upper bound.
-Due to the constraints imposed by legal chess moves, such a fully connected weighted network is not achievable in practice.
-The example shown in Figure 1, which corresponds to the ply with the highest tension, has a tension value of approximately $T\simeq 6.65$.
-For reference, the maximum tension that can be reached by two players iteratively choosing legal moves that most increase tension is around $T\simeq 10.3$ at ply 52, with limited room for further increases (see Supplementary Materials).
-Result 1: AIs tolerate more tension
-We show the typical evolution of the strategic tension throughout many games in Figure 2.
+Small, brownish nodes represent vacant squares under control.
+The chosen ply represents the highest tension reached during the game according to the maximum eigenvalue of the adjacency matrix.
+Among the many ways to characterize the interaction network, we focus on how game complexity evolves, reflected in features such as the number of interacting pieces and loops  [^2] A loop consists of a closed sequence of interactions between pieces forming a cycle within the network (e.g., a white rook on b4 defends a white knight on e4; the knight, in turn, attacks a black bishop on d6, which attacks the rook on b4)..
+We interpret this complexity both in terms of structural richness and strategic fragility, namely the potential for cascading consequences following a suboptimal move.
+In dense, highly connected configurations, even a small error may propagate across the network, shifting advantage substantially.
+To capture this latent vulnerability, we use the maximum eigenvalue of the adjacency matrix, $\lambda_{1}$, adopting the chess-inspired term strategic tension.
+This quantity summarizes how tightly coupled the interaction network is at any moment and reflects critical structural dynamics [31].
+For simplicity, we consider the undirected tension network, treating tension as a reciprocal property between pieces.
+The maximum eigenvalue reflects network functionality, encompassing both its structural features and dynamical properties [8].
+This can be understood via topological entropy, which captures the complexity and unpredictability of paths in a network [1]: the spectral radius summarizes the network’s “diversity” of possible trajectories.
+Higher $\lambda_{1}$ indicates greater dynamical complexity, consistent with tension as a measure of uncertainty or potential instability.
+Importantly, $\lambda_{1}$ is bounded between the average and maximum node degree, $\langle s\rangle\leq\lambda_{1}\leq s_{\max}$, linking global connectivity with local heterogeneity [41].
+Thus, the spectral radius serves as a natural metric for tension, bridging structural properties (as ensured by the Perron–Frobenius theorem) and dynamical behavior (as reflected in entropy-like measures) at the current state of the board (see the Supplementary Materials for details).
+RESULT 1: humans peak, AI sustains STRATEGIC TENSION
+We illustrate the typical evolution of the strategic tension across multiple games in Fig.
+2.
 We can distinguish four main phases.
-The first phase, corresponding roughly to the opening stage up to around ply 30, features a buildup where pieces start approaching each other, and contested squares come under threat, leading to an abrupt increase in tension.
-The second phase, spanning approximately from ply 30 to ply 50, represents the peak of complexity, where the board remains highly interconnected and the tension stays elevated.
-The third exhibits a simplification of the tension network, as exchanges and reductions gradually lower the overall tension to about ply 120.
-The final and fourth is characterized by a low but relatively stable tension, as the remaining pieces maneuver without significantly altering the balance.
-The third and fourth phases are strongly influenced both by the overall game length and by the eventual outcome.
-Thus, the tension captures the overall contours of the game, as discussed in the theory of chess play [10], and indicates that the maximum eigenvalue is a meaningful proxy for game complexity and the potential for cascading effects.
-We do not report tension values beyond ply 150, as the available statistics become increasingly sparse—especially for games played between humans, which typically conclude earlier (Figure 2b).
-We compare the tension profile with the Stockfish evaluation of the same games, which quantifies the game board’s state.
-We find that sharp changes in evaluation often occur at or shortly after peaks in tension [2], supporting the idea that these moments mark critical points where a single suboptimal move can trigger significant consequences (see Supplementary Materials).
+The first phase, corresponding roughly to the opening stage up to around ply 30, features a buildup where pieces approach each other, and contested squares come under threat, leading to an abrupt increase in tension.
+The second phase, spanning approximately from ply 30 to ply 60, represents peak complexity, where the board remains highly interconnected, and the tension stays elevated.
+The third phase exhibits a gradual simplification of the tension network, as piece exchange and loss steadily lower the overall tension.
+Finally, the games reach a relatively stable value of tension, but at this point available statistics become increasingly sparse, especially for games played between humans that typically end earlier (Fig.
+2b and 2c).
+Thus, the dynamics of tension capture the overall contours of the game, including the opening buildup, the combinatorially complex midgame, and the relaxation leading up to the endgame as discussed in the theory of chess play [13].
+Despite the overall similarities in the typical trajectory of strategic tension between human and AI play, there are notable differences.
+Maximum tension is higher in human-played games than in engine-played games.
+Up to approximately ply 40, human players typically follow established opening repertoires from theoretical databases, which often lead to well-studied positions with numerous direct piece interactions, including attacks, defenses, and immediate threats [15, 6].
+These theory-driven positions generate the highest points of tension as players transition to original play in the midgame.
+Once players leave the realm of memorized theory and must rely on their own calculation and evaluation, they tend to simplify positions.
+Human players reduce tension more rapidly and, as we discuss below, resolve piece conflicts sooner, often favoring attack-oriented configurations.
+In contrast, state-of-the-art engines tend to maintain more distributed, positional advantages that unfold gradually over multiple moves, resulting in a slower decay of tension.
+A secondary difference emerges in how strategic tension declines in drawn vs.
+decisive games.
+Drawn games between humans show a rapid and substantial fall in tension relative to decisive games.
+This agrees with the faster decrease in the number of pieces for drawn games between human players (see Fig.
+10 in the Supplementary Materials).
+AI games show a contrasting pattern.
+Drawn games show significantly more (not fewer) pieces on the board relative to decisive games in the endgame and tension does not differentiate strongly.
+II Result 2: Strategic tension reflects, does not predict game result
+To relate the structural complexity of the interaction network to game outcomes, we compare the evolution of strategic tension with the algorithmic evaluation provided by Stockfish (see Fig.
+8 in the Supplementary Materials).
+Inspecting the evaluation score, we identify bifurcation points in game trajectories, typically around ply 30 (move 15), where decisive and drawn games begin to diverge for human players during the opening-to-midgame transition.
+A similar behavior is observed in AI games, where a local maximum or inflection point in the evaluation profile precedes the divergence of outcomes.
+Notably, in all games, this divergence occurs before strategic tension reaches its absolute peak.
+This ordering suggests that the game’s outcome is largely determined during the ascent toward peak tension, before the subsequent reduction in tension differentiates decisive games from draws.
+In chess terms, foundational advantages—such as positional pressure, space, or superior piece coordination—are established while the board is still increasing in complexity.
+The ensuing reduction in tension reflects a phase of conversion or simplification from the tension peak, often marking the resolution of a struggle in which one side has attempted to generate complications.
+Consistent with established principles of high-level play, players with a decisive advantage tend to trade pieces to limit counterplay [37].
+Our results confirm that this reduction is not the cause of the outcome, but rather its consequence: both humans and engines simplify positions once a win or a forced draw is already favored.
 Figure 2: Tension values during games for both humans (grandmasters) and AI (Stockfish and Leela Chess Zero).
@@
 (c) The same analysis is performed for AI games, revealing a similar trend.
-Following the peak, the rate at which the tension decreases depends on several factors.
-Among these, whether a game ends in a draw or not plays only a marginal role compared to the distinction between human-played and engine-played games.
-Generally, draws are significantly more common among chess engines, and tension lasts longer (Figure 2c).
-Moreover, the peak tension typically occurs at a later ply and tends to be lower on average compared to human games.
-Within human games, we observe that those ending in a draw tend to exhibit lower tension.
-This is likely due to many drawn games’ more cautious and symmetrical nature, where early simplifications and fewer sharp imbalances keep the tension low.
-In contrast, tension values are not as clearly differentiated between drawn and decisive outcomes in games played between AIs.
-Result 2: Complementary Metrics Beyond Tension
-Tension captures general and coarse features of the game.
-For example, AI maintains a higher number of pieces on average, reflecting less simplification early on.
-The interaction networks derived from engine games display more links and loops, indicating a richer and more resilient web of mutual threats, defenses, and controls.
-That the maximum average number of loops is, on average, higher in human games, as with the higher peak in tension, indicates similarity between the two metrics, which are discussed further in “ Supplementary Materials.”
+RESULT 3: Variability in strategic tension reflects different gameplay
+While tension seemingly captures only general and coarse features of the game, it is correlated with key aspects of gameplay that distinguish humans from AI (see Fig.
+9 in the Supplementary Materials).
+Number of links and loops: AI interaction networks exhibit more links and loops, indicating a richer web of mutual threats, defenses, and controls.
+However, the maximum average number of loops is higher in human games, reflecting the peak of human tension.
+Number of pieces: AI maintains a higher number of pieces over the trajectory of games on average.
+This reflects a weaker tendency to simplify the game by removing pieces from the board.
+From mid to endgame, the small number of pieces limits the strategic tension that is typically reached in human games.
 Figure 3: Comparison of structural properties across human and AI games.
-We analyze the set of games shown in Figure 2.
+We analyze the set of games shown in Fig.
+2.
 Two measures are considered: the average standard deviation of the degrees associated with chess pieces, normalized by the mean of the degrees, and the average tension per piece.
 Red is for AI and blue for human games.
-As with tension, engine play in the endgame shows more variation in how pieces are connected to one another.
-We measure this with the noise coefficient of the degree distribution over chess pieces (standard deviation over the mean degree), which follows a similar trend in both human and AI games up until the final stages, where they start to diverge at approximately ply 80 (Figure 3a).
-There, AI exhibits greater variability in the degree distribution.
-This indicates that, on average, human players tend to distribute interactions more evenly across their pieces, while AI tends to stress interactions on specific pieces.
-This pattern aligns with the higher tension values in the endgame because a more heterogeneous distribution of degrees facilitates the emergence of highly connected hubs, increasing the adjacency matrix’s maximum eigenvalue.
-This difference in tension between AI and human play, however, is inverted when normalized by the number of chess pieces remaining on the board, $\lambda_{1}/n_{t}$.
-As shown in Figure 3b, it remains consistently higher across all plies in human games.
+Degree distribution: The noise coefficient of the degree distribution over chess pieces (defined as the standard deviation over the mean degree) mirrors the rise and fall of strategic tension, first falling then rising.
+While human and AI games display a broadly similar global trend, human play exhibits more pronounced variation as shown in Fig.
+3a.
+In particular, during the opening phase, the relative degree variation decreases markedly, reflecting a positioning phase in which interactions are more evenly distributed across pieces as tension accumulates.
+As the game progresses, however, the interaction network in human games exhibits greater variability, eventually returning to levels similar to those observed in AI games, or even lower.
+Tension per piece: As shown in Fig.
+3b, the tension per piece $\lambda_{1}/n_{t}$ remains consistently higher across all plies in human games.
 While human play shows an approximately linear increase in tension per piece throughout the game, AI games display a sublinear progression that appears to plateau around ply 80.
-This suggests that while AI maintains “tactical pressure” more persistently during late-game phases compared to human play, it does so with significantly more pieces remaining on the board.
-To quantify better this effect, we also investigated the correlation between tension values and the number of chess pieces present on the board (see Supplementary Materials).
-While we have thus far treated edges in the graph equally, they are different.
-In particular, the surplus of attack links over defense links, or Attack-Defense Balance (Figure 4), highlights how the early phase of the game tends to be more defense-oriented.
-As the game progresses, however, attack links gradually outnumber defense links.
-This shift is noticeably more pronounced in human games compared to those played by AI.
+Human games involve more frequent captures of pieces (see Supplementary Materials), but the remaining pieces are then involved in a greater number and greater complexity of interactions.
+Attack-defense balance: When distinguishing between attack and defense links, we find that the surplus of attack links over defense links, or “attack-defense balance” as shown in Fig.
+4, moves inversely with the tension profile.
+It first drops in the early phase more defense-oriented phase of the game.
+As the game progresses, attack links gradually outnumber defense links.
+The curve is shifted towards more positive values in human games compared to AI games.
 Figure 4: Balance between attack and defense links across human and AI games.
-We analyze the set of games shown in Figure 2.
+We analyze the set of games shown in Fig.
+2.
 For both human and AI games, we distinguish between games that end in a draw and those that result in a win/loss, represented with light and dark shades, respectively.
-This behavior depends on different strategic styles, especially for drawn AI games, which a lower Attack-Defense Balance characterizes.
-Human players often seek direct and decisive operations, favoring sharper, more aggressive transitions when tactical chances arise.
-In contrast, AI maintains a more balanced configuration in the long run, resulting in wider piece coordination, latent threats, and structurally embedded defenses.
-In draws, AI games prefer solidity over risk, with extended defensive networks and fewer explicit attacks.
-This pattern also aligns with the tension trends shown in Figure 2, where tension remains high and decreases slowly in drawn AI games.
-Since defense links can only form in response to attacks, a high tension level in the later phases of the game likely reflects a dense but balanced interaction, where attacks tend to be answered by appropriate defenses.
-Extending the parallel with the tension profile, we observe that in this case, it is human games that reach a higher peak of tension, whereas in the case of the Attack-Defense Balance, AI games exhibit a more pronounced minimum.
-This comparison shows that tension aligns with a balance between attack and defense.
-Overall, we find that strategic tension aligns with basic metrics about gameplay, such as the number of pieces and loops in the interaction graph.
-Still, it is likewise helpful to compare it with the balance across pieces on the board and how the pieces are arranged in formation.
-The picture is that AIs maintain more pieces, a more balanced network of attack and defense connections, and display more variability in the tension across the pieces they retain.
-Result 3: Tension increases with skill
+Overall, we find that strategic tension aligns with basic gameplay metrics, such as the number of pieces and, in particular, the presence of loops in the interaction graph.
+At the same time, it is informative to relate it to the balance across pieces on the board and to how pieces are arranged in formations.
+Taken together, our results indicate that AI games retain more pieces, exhibit a more balanced network of attack–defense connections, and display less variability in the degree distribution across pieces.
+RESULT 4: Cumulative TENSION SCALES WITH SKILL AND COMPUTATIONAL COMPLEXITY
 Figure 5: Tension loads for different Elo ratings and Stockfish depths.
-(a) We represent tension loads (sum of tension values up to ply 150) for human games (blue dots) and AI games between Stockfish and Leela from official TCEC tournaments (red dot), all associated with estimated Elo ratings.
-Each dot represents the average across 1,200 games.
-The blue dot associated with the highest Elo rating and the red dot were derived from the data analyzed in Figure 2.
+(a) We represent tension loads (sum of tension values up to ply 150) for rapid games (turquoise dots), blitz games (green dots), classical games played by top human players (blue star), and games between Stockfish and Leela from official TCEC tournaments (red star), all associated with estimated Elo ratings.
+Each dot/star represents the average across 1,200 games.
+The blue and red stars were derived from the data analyzed in Fig.
+2.
+For reference, blitz games are played with very short time controls (typically 3–5 minutes per player), rapid games with intermediate time controls (around 10–25 minutes per player), and classical games with long time controls (often 60 minutes or more per player) allowing deeper strategic play.
 (b) We represent tension loads from games between versions of Stockfish at varying depth levels (orange dots), computed from a smaller sample of 120 games per dot.
-While engines can maintain higher tension for a longer duration compared to humans, how does this difference depend on skill level?
-We consider the cumulative tension up to ply 150 across different Elo ratings for human players in Figure 5a  [^3] In principle, Stockfish has a search depth parameter that serves as a proxy for skill level.
-Leela has a time limit parameter, during which it is permitted to assess the state of the board and make a decision.
+The two panels also show the $R^{2}$ values of the corresponding linear regressions.
+While engines can maintain higher tension for a longer duration compared to humans, does this difference depend on skill level?
+We consider the cumulative tension up to ply 150 across different Elo ratings for human players in Fig.
+5a  [^3] In principle, Stockfish has a search depth parameter that serves as a proxy for skill level.
+Leela has a time limit parameter during which it is permitted to assess the board state and make a decision.
 At a few dozen search depths, the tension load is limited to about 450, similarly for Leela at a few seconds.
 Since the games that are shown go far beyond these limited parameter values with supercomputers, this suggests that algorithmic complexity increases the ability to handle tension..
-Tension changes substantially with Elo, revealing discrete jumps that separate human players into roughly three, significantly separate groups.
-A plateau in tension load from Elo of 1000 ends between 1500 and 1700.
-This transition typically corresponds to the entry point into competitive chess, where players begin to study openings, apply strategic principles more consistently, and reduce the frequency of major tactical blunders.
-After a second plateau, a jump in tension occurs at 2100 Elo.
-This point marks the entry into the realm of stronger players, such as experts and grandmasters, whose games exhibit more nuanced, complex positions with deeper strategic and tactical layers.
-Figure 5b shows the tension load values for various fixed search depths of Stockfish.
-These results indicate an approximately constant tension load across different depths, suggesting a qualitatively different scenario compared to the Stockfish vs.
-Leela games from the TCEC tournaments (highlighted as the red point in Figure 5a), where the engines are free to adjust their search depth and, more generally, the time allocated per move  [^4] Representing the dependence of tension on the time limit per move in the case of Leela is challenging because it strongly depends on the technical specifications of the machine used, particularly the presence of a suitable dedicated GPU and sufficient RAM.
-In particular, achieving Elo ratings significantly higher than the best human players requires substantial hardware resources..
-The correlation between Stockfish search depth and Elo rating is not well established.
-Previous experiments on a different engine (Houdini) reported a roughly linear relationship up to a search depth of 20 plies [13].
-II Discussion
+Tension changes substantially with Elo, showing an approximately linear dependence in human games.
+Moreover, across all Elo values, rapid games consistently exhibit slightly higher tension loads than blitz games.
+This is consistent with the idea that longer time controls allow human players to create and manage more complex positions [5].
+The effect is even more pronounced in classical games, where players have significantly more time to make decisions, resulting in higher tension and reflecting the chess community’s concern that time-constrained players experience larger performance drops on complex decisions and therefore tend to choose safer, less strategically risky continuations [5, 7].
+In contrast, elite classical play is often characterized by keeping simplification at bay and maintaining a complex struggle, where grandmasters sustain tense, fluid positions that test the opponent’s stamina and depth of calculation over many hours [22, 30, 12].
+For comparison, we examine how the tension load changes for Stockfish, where we can control the computational complexity of the algorithm by adjusting the search depth at a fixed value for simulated games in Fig.
+5b.
+Although we examined search depths only up to 50, the results already indicate that increasing depth leads to higher tension.
+As a reference, previous experiments using a different engine (Houdini) reported a roughly linear relationship between Elo rating and search depth up to 20 plies [17].
+Overall, this indicates that increased tension load is indicative of greater computational demands on the player.
+III Discussion
 We investigate the dynamics of strategic tension, a network-based definition of piece interactions, in chess games played by humans and AI.
 Our central hypothesis is that distinct tension profiles would emerge, reflecting differing strategic approaches.
-The results confirm this hypothesis and reveal several key insights.
-First, we observed that state-of-the-art chess engines (Stockfish at high search depths and Leela Chess Zero at advanced neural network settings) consistently maintain significantly higher levels of strategic tension for longer durations than even the best grandmasters.
-This suggests an ability to navigate and sustain complex, balanced positions where multiple pieces exert pressure and threats remain unresolved.
-Second, human players exhibit a notable relationship between skill level (Elo rating) and tension management.
-Significant differentiation in accumulated tension emerges primarily above an Elo rating of approximately 1600, coinciding with the stage where players typically develop deeper strategic understanding, theoretical knowledge, and enhanced calculation abilities [9, 7].
-This implies that managing higher levels of strategic tension is characteristic of advancing chess expertise in humans.
-Likewise, one might expect greater computational complexity to allow an engine to find and sustain even more complex positions, leading to a monotonic increase in tension load with search depth, similar to the trend observed with human expertise.
-Preliminary calculations indicate that the computational complexity of Stockfish, as reflected in its search depth, does influence tension load, but in a nonlinear manner: tension load exhibits flat plateaus across a wide range of fixed search depths (approximately from 10 to 30).
-This raises the question of how and where new play styles emerge as the complexity accessible to the algorithm increases.
-The observed differences in tension dynamics distinguish how biological and artificial systems approach problem-solving under strategic constraints.
-Human cognition appears adapted to manage, and perhaps even actively reduce, high levels of sustained complexity, possibly due to inherent limitations in working memory, attention, or susceptibility to psychological pressure [15, 18].
-State-of-the-art AI, optimized for performance within defined rule sets, demonstrates a capacity for tolerating and exploiting complexity [30].
-The superior ability of high-level engines to sustain tension elucidates their ability to handle immense combinatorial complexity [34, 22]: engines do not simply calculate better but actively maintain more complex strategic states.
-This aligns with the idea that optimal strategy might be contingent on the agent’s cognitive or computational architecture.
-More broadly, such differences may be essential for AI systems designed for human interaction.
-While direct extrapolation requires caution, we might think of how to examine tension in other complex adaptive systems, such as economic markets or geopolitical negotiations [32].
-The finding that increased computational complexity allows AI to sustain higher tension may matter for stability in international relations, potentially enabling longer standoffs but also increasing the stakes should tension break [14, 31] in an AI-enabled analogy to Dr.
+Our observations confirm this hypothesis and reveal that state-of-the-art chess engines (Stockfish at high search depths and Leela Chess Zero at advanced neural network settings) consistently maintain significantly higher levels of strategic tension for longer durations than even the best human grandmasters.
+We speculate that the sustained tension in AI play reflects the deferral of immediate opportunities and facilitates latent strategic options that only materialize in later phases of the game.
+This kind of strategic decision making would be facilitated by the ability to handle greater computational complexity.
+Correspondingly, we find that human players exhibit a linear relationship between cumulative tension and skill level (Elo rating).
+Similarly, Stockfish’s tension load increases with computational complexity, as reflected in its search depth.
+Far beyond these examples are AI engines running on supercomputers.
+Superhuman computational capacity obviates the need to simplify upon leaving theoretical territory; AI engines do not need to rely on heuristics and intuition such as familiarity with positions in opening theory that heavily guide expert human play.
+These observations suggest that the ability to sustain strategic tension is a hallmark of (for human players) advanced chess expertise, encompassing deeper strategic understanding, and extensive theoretical knowledge and (for AI) enhanced calculation skills [12, 9].
+The increase in tension with computational load likely arises from different factors in the two groups, reflecting their nature as artificial versus biological systems.
+In AI, tension rises as engines gain computational capacity and play at higher levels, where controlling more pieces and squares maximizes options and exploits opponents’ mistakes [4].
+In humans, more skilled players can handle complex positions and maintain higher‑tension situations without simplifying [23, 11, 10, 18].
+Indeed, humans may tend to simplify under certain conditions possibly due to limitations in working memory, attention, or susceptibility to psychological pressure [20, 26].
+This interpretation of the results could also suggest that human players may sometimes actively steer games toward drawn resolutions, reflecting deliberate risk minimization and metagame considerations typically absent in AI play.
+The simplification of high-tension positions thus serves not only strategic purposes but also as an adaptive response to the computational demands of accurate tactical calculation, allowing players to sustain the level of accuracy needed for strong play within cognitive limits [14].
+Of course, such limits come from the evolutionary trade-offs reflected in the brain, where short-term costs, such as losing a chess game, are outweighed by long-term benefits in metabolic efficiency and general, rather than domain-specific, intelligence [snell-roodBrainSizeGlobal2009, burgerAllometryBrainSize2019].
+Several avenues for future research emerge from this work.
+Extending this methodology to other complex strategic games, such as Go [38], Shogi, or even real-time strategy e-sports, could test the universality of the observed dynamics across different game structures (e.g., imperfect information and varying spatial dynamics).
+Moreover, the AI engine used (Stockfish) represents one architecture; other engines with different algorithms, such as neural network-based engines like AlphaZero, may show different tension dynamics.
+While we focused on Stockfish and Leela due to their prominence and availability, our analytical framework and open-source code can readily be applied to other UCI-compatible engines, enabling broader comparative studies.
+With a more direct focus on tension, directly investigating the cognitive and physiological correlates of our tension metric in human players—for example, by combining game analysis with eye-tracking, think-aloud protocols, or stress indicators such as galvanic skin response—could help elucidate the underlying mechanisms driving changes in tension.
+The observed differences between artificial and human systems could have implications for AI designed to interact with humans, particularly as AI is expected to play an increasing role in teaching and training across many domains.
+In chess, a notable step in this direction is the Maia engine [29], which is explicitly designed to model human play rather than engine-optimal moves.
+Moreover, while direct extrapolation requires caution, examining how tension is sustained in other complex adaptive systems, such as economic markets or geopolitical negotiations, can provide insight [35].
+The finding that increased computational complexity leads AI to sustain tension may matter for stability in international relations [simmons-edlerPositionAIPowered2024], potentially enabling more complex standoffs but also increasing the stakes of miscalculation [19, guttieriIntegrativeComplexity1995, 34], creating an AI-enabled parallel to Dr.
 Strangelove.
-Several avenues for future research emerge from this work.
-The network-based tension metric represents one specific operationalization of strategic complexity.
-It primarily captures piece interactions and may not fully encompass other crucial factors like pawn structure, king safety, or long-term positional plans.
-While it changes predictably with several measures of board complexity (Result 2), parallel measures may help inform pedagogical approaches in chess.
-Further investigations could also focus on how different engines manage tension, examining variations across engine architectures, versions, and underlying algorithms.
-Extending this methodology to other complex strategic games, such as Go [36], Shogi, or even real-time strategy e-sports, could test the universality of the observed dynamics across different game structures (e.g., imperfect information, different spatial dynamics).
-Directly investigating the cognitive and physiological correlates of our tension metric in human players—for example, by combining game analysis with eye-tracking, think-aloud protocols, or stress indicators (like galvanic skin response)—could help elucidate the underlying mechanisms driving human tension.
-Furthermore, exploring whether AI can be explicitly trained to modulate a tension profile, perhaps to mimic human playstyles or achieve specific pedagogical goals, represents an interesting direction for human-AI interaction research.
-Investigating the relationship between our tension metric and game outcomes (win/loss/draw) beyond simple averages could also reveal critical thresholds or patterns predictive of game resolution.
-Finally, our analysis relies on existing game databases.
-While large, these datasets may have inherent biases, and game conditions (e.g., time controls, stakes) might differ from controlled experimental settings.
-The specific AI engine used (Stockfish) represents one architecture; other engines with different algorithms (e.g., neural network-based like AlphaZero beyond its initial papers) might exhibit different tension dynamics.
-While this study focused on Stockfish and Leela due to their prominence and availability, the analytical framework and open-source code we provide can be readily applied to other UCI-compatible engines, paving the way for broader comparative studies.
-We find evidence for distinct approaches to managing strategic tension between human players and AI chess engines.
-By proposing a network-based metric, we characterize how tension typically accumulates and dissipates throughout games, revealing that high-performance AI sustains significantly greater complexity for longer periods than even elite human players.
-Human tension management appears correlated with expertise but potentially constrained by cognitive factors, while AI’s capacity in tension management is primarily a function of computational resources, which allows it to consider high-dimensional configuration of balanced interactions.
-While human players would find such algorithmic complexity inaccessible, the brain is the outcome of a trade-off in terms of longer-term evolutionary games.
-The short-term cost of losing a chess game might be worth the long-term benefits of metabolic cost and cognitive plasticity in the game of life.
 Code availability
@@
 Authors’ contributions
-AC developed the code, performed the numerical analyses, and generated the figures.
-VDPS provided the original inspiration for the project and supervised its development at all stages.
-EDL participated in discussion.
-All authors contributed to writing the manuscript.
+AC led the definition of the tension metric, with contributions from the other authors, and carried out the code development, numerical analyses, and figure generation.
+EDL performed several simulations on a computing cluster using Stockfish at high depths.
+VDPS provided the original inspiration for the project, supervised its development at all stages, and independently cross-checked the results.
+All authors participated in the discussions and contributed to writing the manuscript.
 Acknowledgements
@@
 It was also partially funded by the Austrian Science Fund (FWF) 10.55776/ESP127.
-III Methods
-III.1 Data
+IV Methods
+IV.1 Data
 Chess games are commonly stored in the PGN (Portable Game Notation) format.
@@
 The typical time control is 120 minutes per player, with a 12-second increment added after each move.
-These datasets served as the basis for the analyses presented in Figure 2.
-In addition, a large set of games played by human players with Elo ratings ranging from 1100 to 2300 was downloaded from https://www.ficsgames.org/, an online archive that collects games played on the Free Internet Chess Server (FICS).
-This dataset broadly represents amateur to intermediate-level players, allowing for analysis across a wide skill spectrum.
-Together with the previously described datasets, these data formed the basis for the analyses presented in Figure 5.
-III.2 Software
+These datasets served as the basis for the analyses presented in Fig. 2.
+In addition, a large set of rapid and blitz games played by human players with Elo ratings ranging from 1000 to 2700 was downloaded from https://lichess.org/, an online platform that provides a vast archive of games.
+This dataset broadly represents amateur to expert players and is ideal for analysis because games are subject to anti-cheating controls, ensuring reliable data across a broad skill spectrum.
+Together with the previously described datasets, these data formed the basis for the analyses presented in Fig. 5.
+IV.2 Software
 All operations related to chess move generation, rule validation, and position analysis were carried out using the python-chess library (v1.9.4).
@@
 Furthermore, it provides native support for communicating with UCI-compatible engines such as Stockfish, which allowed us to automate the evaluation of positions and simulate engine-vs-engine games with varying search depths.
-III.3 Generating Stockfish games at varying depths
+IV.3 Generating Stockfish games at varying depths
 To generate the games between Stockfish and itself at various search depths, we followed a procedure designed to balance realism and controlled variability.
@@
 This asymmetry introduced slight variability in evaluation and decision-making while preserving a consistent overall depth level $D$.
-IV Supplementary materials
-IV.1 Why the spectral radius as a measure of tension
-Topological entropy, $h_{\mathrm{top}}$, is a concept from dynamical systems theory that quantifies the exponential growth rate of the number of distinguishable orbits or, in the context of graphs, the diversity of paths.
-For unweighted networks (where the adjacency matrix has entries in $\{0,1\}$), the spectral radius $\lambda_{1}$ of the adjacency matrix is directly related to topological entropy.
-In particular, it has been shown that $\lambda_{1}$ satisfies a variational principle equivalent to the dynamical entropy of the network (i.e., the entropy of a stochastic process on the graph), leading to the definition of topological entropy as
-$h_{\mathrm{top}}=\log(\lambda_{1}).$
-Topological entropy was first introduced in 1965 by Adler, Konheim, and McAndrew [1] as a measure of the complexity of a dynamical system.
-In the context of networks, this relationship is particularly meaningful: the quantity $\lambda_{1}$ does not merely reflect the number of walks in the graph, but its logarithm measures the average Shannon information per step gained when observing a random walk unfold.
-A larger $\lambda_{1}$ indicates a greater number of possible distinct paths at each step, implying higher unpredictability, richer structural diversity, and overall complexity.
-This connection positions $\lambda_{1}$ as a fundamental quantity bridging structural graph theory, information theory, and dynamics.
-It can be interpreted as a proxy for the “intrinsic computational capacity” or “information-processing potential” of the network.
-In our study, this interpretation is particularly useful: we use the spectral radius $\lambda_{1}$ as a global measure of the tension and complexity present on the chessboard.
-A higher $\lambda_{1}$ suggests a denser and more interconnected pattern of potential interactions between pieces (e.g., threats and defenses), and therefore a richer position in tactical possibilities and strategic depth.
-This justifies using $\lambda_{1}$ as a meaningful descriptor of the game’s complexity at a given state.
-IV.2 Maximized tension game
-If we consider the unrealistic and absurd game in which, at each ply, the players randomly choose one of the moves that maximizes tension, by ply 52 the tension reaches a value of $T\simeq 10.3$ (Figure 6).
-Figure 6: Tension network obtained by maximizing tension move by move.
-(a) Chessboard at ply 44, obtained by choosing the legal move that increases the tension the most at each ply.
-(b) The corresponding network represents the interactions on the chessboard, using the same color and edge scheme as in Figure 1.
-The first 60 moves of the game are listed below: 1.
-d3 h6 2.
-Bg5 e5 3.
-Be7 Nc6 4.
-Qd2 d5 5.
-Qxh6 Qd6 6.
-Qe6 Nd8 7.
-c4 Rh6 8.
-cxd5 e4 9.
-Nc3 g5 10.
-Na4 b5 11.
-Rc1 Rb8 12.
-Rxc7 Rb7 13.
-Nc5 Qxd5 14.
-dxe4 Rb6 15.
-e3 Ra6 16.
-Rxa7 b4 17.
-Bd3 Bg7 18.
-Nf3 Bf8 19.
-Nxg5 Bd7 20.
-f4 Bc8 21.
-Kf2 b3 22.
-Rc1 bxa2 23.
-Rc3 a1=Q 24.
-Ra3 Qd1 25.
-Ra4 Qg4 26.
-Be2 Qgf5 27.
-Bd3 Qfxe4 28.
-Ra8 Nxe7 29.
-Kg3 Rh4 30.
-Kf2 Rh6.
-After ply 52, the game maintains a relatively constant level of tension indefinitely, indicating that the game ceases to progress meaningfully.
-IV.3 Stockfish evaluation for humans and engines
-To better analyze the progression of games, we also examined the Stockfish evaluation, again distinguishing between games played by humans or AIs, and whether they ended in a draw or not (Figure 7).
-What emerges is that games between AIs show a rapid increase in the absolute value of the Stockfish evaluation from the very first plies.
-As the game progresses, the average evaluation values begin to diverge depending on the eventual outcome: games that end in a win or loss exhibit a steady increase in evaluation with each ply, whereas those that result in a draw maintain relatively low evaluation values throughout.
-In contrast, in human games, the evaluation remains near zero during the opening phase.
-It is worth noting that, in fact, only starting from 2023 did official rules begin allowing Leela Chess Zero to select its openings directly from a database.
-However, as the game progresses, a similar divergence in evaluation can be observed.
-Notably, this shift typically occurs around ply 40, where tension generally peaks, marking a transition where evaluations for decisive games begin to rise consistently.
-In contrast, those for drawn games remain comparatively stable.
-Barthelemy [2] has already observed this result, using a similar framework for capturing the notion of fragility within chess games.
-In the case of human games, the average values of tension are characterized by greater fluctuations, likely due to the lower predictability of human play and, more specifically, the greater variability between individual games, especially when different players are involved.
-Moreover, when focusing on tension values associated with drawn games, a notable difference emerges: in the case of AI matches, tension decreases towards zero much more gradually.
-This is mainly because engine games are longer on average than human games.
-Figure 7: Stockfish evaluation during games for humans and AI.
-We consider 120 games for each (a subset of those analyzed in Figure 2), with evaluations computed at a fixed depth of 20.
-The same color scheme described in Figure 2 is used: cold and warm colors for humans and AI, respectively, and light/dark shades to distinguish draws from decisive games.
-We also show the overall average evaluations, along with vertical dashed lines indicating the ply corresponding to the maximum tension for both groups, humans and AI, shown in blue and red, respectively.
-To better understand the type of information conveyed by the Stockfish evaluation, we analyzed the 2,400 games to compute the probability that a player goes on to win the game when holding an evaluation advantage of at least 100 centipawns at ply 40.
-In the case of human games, this probability is approximately 71 %, whereas for AI games, it is around 59 %.
-IV.4 Additional properties of the chessboard and network
-In this section, we present further comparative metrics between human and AI chess games, including the number of pieces, the sum of the weighted material in the chess board - where pieces are valued at 1 (pawns), 3 (knights and bishops), 5 (rooks), 9 (queens), and 0 (kings) - along with the number of links and the number of loops in the tension network.
-These measures provide complementary insights into positional advantages and dynamic board conflicts across different types of games (Figure 8).
-IV.5 Correlation between tension and the number of chess pieces
-We investigated the correlation between the measured tension values and the number of chess pieces present on the board.
-Naturally, this correlation is highly dependent on the stage of the game being considered.
-For this reason, we analyzed three distinct moments in the game—specifically at ply 30, 60, and 90—to show how, on average, tension varies with the number of pieces on the board, both for the AI and for human players (Figure 9).
-We observe that, on average, tension increases with the number of chess pieces across all three stages of the game.
-However, this increase diminishes rapidly as the number of pieces grows, eventually reaching an approximately constant behavior—or even a slight decrease.
-This may be because when the number of pieces is significantly higher than what would typically be expected at a given ply, it might indicate that no major confrontation has yet occurred.
-Such confrontations often lead to a rise in tension, as more pieces are brought into key positions on the board.
-Moreover, an excessive number of pieces on the board can severely limit their mobility, reducing their ability to control important areas or pose threats to the opponent.
-Therefore, it is reasonable to expect that, to maximize tension, there exists an ideal range for the number of pieces — neither too high nor too low.
-Figure 8: Comparison of additional structural properties across human and AI games.
-We analyze the set of games shown in Figure 2.
-Four measures are considered: the number of pieces, the weighted material, the number of links, and the number of loops in the interaction network.
-The color scheme is red and blue for AI and human games, respectively.
-Figure 9: Analysis of the dependence of tension on the number of chess pieces at three game moments.
-We analyze the set of games shown in Figure 2.
-Average tension values are shown for each number of chess pieces, distinguishing between three game moments: ply 30, 60, and 90.
-The color scheme is red and blue for AI and human games, respectively.
-The error bars correspond to the standard deviation calculated for each represented value.

### v3

@@
 In the context of an individual’s lifetime, these can be characterized as strategic choices that involve an implicit trade-off between overfitting the local environment and maintaining the flexibility to adapt to unforeseen environmental conditions [27], in an echo of the speed-accuracy trade-off [33].
-In many real-world settings, a complete measurement of the strategic environment, costs, and outcomes is not available.
-Chess, by contrast, is a game of perfect information, providing a well-studied and bounded setting.
+In many real-world settings, complete measurements of the strategic environment, costs, and outcomes are not available.
+Chess, by contrast, is a game of perfect information, offering a well-studied, bounded setting.
 As a result, chess has served as a model system for investigating strategy.
-It has well-defined rules, perfect information, complexity, and rich history as a testbed for both human expertise and artificial intelligence (AI) [36, 39, 11, 32].
+It has well-defined rules, complexity, and rich history as a testbed for both human expertise and artificial intelligence (AI) [36, 39, 11, 32].
 Chess is a competitive game between two players, white and black, who take turns making moves with a single piece on a board with 8x8 squares as shown in Fig.
@@
 Control links, finally, represent subtler forms of influence, where empty squares become contested zones.
-In constructing the network, we disregard which side is to move, focusing instead on the full set of interactions at each ply to capture the full state of the board.
+In constructing the network, we disregard which side moves, focusing instead on the full set of interactions at each ply to capture the board’s full state.
 Attack and defense links can have weight 1 or 2, depending on whether the corresponding conditions are satisfied by one or both of the pieces involved.
@@
 (b) The network represents the interactions on the chessboard.
-The chess pieces are depicted as dark and light circles, depending on whether they are black or white.
-Attack links are displayed in orange when originating from white pieces and in red when originating from black.
+The chess pieces are depicted as dark or light circles, depending on their color.
+The attack links are shown in red when originating from black and in orange when originating from white.
 Defense links are turquoise when white is defending and blue otherwise.
@@
 We can distinguish four main phases.
-The first phase, corresponding roughly to the opening stage up to around ply 30, features a buildup where pieces approach each other, and contested squares come under threat, leading to an abrupt increase in tension.
+The first phase, roughly corresponding to the opening stage up to around ply 30, features a buildup in which pieces approach each other and contested squares come under threat, leading to an abrupt increase in tension.
 The second phase, spanning approximately from ply 30 to ply 60, represents peak complexity, where the board remains highly interconnected, and the tension stays elevated.
@@
 Human players reduce tension more rapidly and, as we discuss below, resolve piece conflicts sooner, often favoring attack-oriented configurations.
-In contrast, state-of-the-art engines tend to maintain more distributed, positional advantages that unfold gradually over multiple moves, resulting in a slower decay of tension.
+In contrast, state-of-the-art engines tend to maintain more distributed positional advantages that unfold gradually over multiple moves, leading to slower tension decay.
 A secondary difference emerges in how strategic tension declines in drawn vs.
@@
 This ordering suggests that the game’s outcome is largely determined during the ascent toward peak tension, before the subsequent reduction in tension differentiates decisive games from draws.
-In chess terms, foundational advantages—such as positional pressure, space, or superior piece coordination—are established while the board is still increasing in complexity.
+In chess terms, foundational advantages—such as positional pressure, space, or superior piece coordination—are established as the board becomes increasingly complex.
 The ensuing reduction in tension reflects a phase of conversion or simplification from the tension peak, often marking the resolution of a struggle in which one side has attempted to generate complications.
@@
 4, moves inversely with the tension profile.
-It first drops in the early phase more defense-oriented phase of the game.
+It initially decreases during the early, more defense-oriented phase of the game.
 As the game progresses, attack links gradually outnumber defense links.
@@
 For reference, blitz games are played with very short time controls (typically 3–5 minutes per player), rapid games with intermediate time controls (around 10–25 minutes per player), and classical games with long time controls (often 60 minutes or more per player) allowing deeper strategic play.
-(b) We represent tension loads from games between versions of Stockfish at varying depth levels (orange dots), computed from a smaller sample of 120 games per dot.
+(b) We represent tension loads from games between versions of Stockfish at fixed depth levels (orange dots), computed from a smaller sample of 120 games per dot.
 The two panels also show the $R^{2}$ values of the corresponding linear regressions.
 While engines can maintain higher tension for a longer duration compared to humans, does this difference depend on skill level?
-We consider the cumulative tension up to ply 150 across different Elo ratings for human players in Fig.
+We consider the cumulative tension up to ply 150 (Tension load) across different Elo ratings for human players in Fig.
 5a  [^3] In principle, Stockfish has a search depth parameter that serves as a proxy for skill level.
@@
 AC led the definition of the tension metric, with contributions from the other authors, and carried out the code development, numerical analyses, and figure generation.
-EDL performed several simulations on a computing cluster using Stockfish at high depths.
+EDL performed simulations on a computing cluster using Stockfish at high depths.
 VDPS provided the original inspiration for the project, supervised its development at all stages, and independently cross-checked the results.
@@
 Barthelemy for useful discussions and for sharing his code, which we used as a baseline for our analysis.
-This work was partly funded by the City of Vienna, Municipal Department 7, and the Federal Ministry of the Republic of Austria for Climate Action, Environment, Energy, Mobility, Innovation, and Technology as part of the project GZ2021-0.664.668.
+This work was partly funded by the City of Vienna, Municipal Department 7, and the Federal Ministry of the Republic of Austria for Climate Action, Environment, Energy, Mobility, Innovation, and Technology as part of the project GZ 2023-0.841.266.
 It was also partially funded by the Austrian Science Fund (FWF) 10.55776/ESP127.